OpenStudy (anonymous):

I'm trying to find the argument of the following function but don't understand what i do wrong.

OpenStudy (anonymous):

\[\arg(\frac{ 4.17 }{ (1+ 1.6i)*1.6i}*e^{0.1s})=\arg(4.17\times e^{0.1s})-\arg((1+ 1.6i)*1.6i))\]

OpenStudy (anonymous):

\[=\arg(e^{0.1i})-\arctan(\frac{ 1 }{ -1.6 })\] But I get the right result only when I subtract with pi?